OpenStudy (anonymous):

(x-2)/3 = 2x + 8/6x help..? Please? step by step?

OpenStudy (anonymous):

hello imran

OpenStudy (anonymous):

Hi! (:

OpenStudy (anonymous):

\[\frac{x-2}{3}=\frac{2x+8}{6x} \iff 6x(x-2)=3(2x+8)\] is a start

OpenStudy (anonymous):

multiply out, get a quadratic equation so solve namely \[6x^2-12x=6x+24\] \[6x^2-18x-24=0\] \[x^2-3x-4=0\] and i bet you are good from there because this factors easily

OpenStudy (anonymous):

Thank you very much. I kind of suck at math and have a lot of math homework to do.

OpenStudy (anonymous):

so is it x=4 and x= -1 ?